# CAP[Uncaptioned image] : Evaluation of Persuasive and Creative Image Generation

### v1

@@ 3.4 Persuasiveness
We set two constraints for persuasiveness in the advertisement image generation task: 1. Persuasiveness 2. Alignment with the $AR_{m}$. To address these two constraints, we introduce the Persuasiveness Alignment (PA) score. Because of the complexity of the direct evaluation of persuasion, we decompose the persuasion into the persuasiveness components in [36, 43] and creativity components in [40]. We define a set of questions, which we use to score the persuasion components via a LLM, and to obtain human annotations. The questions are:
@@
  • Audience (AU): Each advertisement is supposed to target specific groups of audiences [43]. How well does the generated image target the audience the message targets?
@@
  • Benefit (B): One way to make the ads more persuasive is to picture the benefits of the product features for the customer. How well does the image convert the features to benefits for customers?
@@
  • Appeal Category (AP): Each advertisement is intended to target one of three appeal categories: Ethos, Pathos, or Logos [36, 4]. Ethos appeals to ethics and credibility; for example, when a celebrity endorses a product, the advertisement targets Ethos. Logos appeals to logic, using reasoning or factual information to persuade. Pathos appeals to emotion, aiming to evoke an emotional response in the audience. How well does the image target the appeal category of the advertisement message?
  • Elaboration (E): How visually detailed is the image [40]?
  • Originality (O): How out of the ordinary and unique is the image [40]?
@@
  • Imagination (I): How well does the image help the audience imagine something they never experienced [40]?
@@
  • Synthesis (S): How well does the image connect elements that are usually unrelated [40]?
To evaluate the persuasiveness of the images, we first generate a description of the image using an MLLM. Next, we utilize an LLM to detect the targeted audience, and the appeal category for the advertisement message. Then, because of superiority of LLMs in reasoning [52, 3, 27], we prompt an LLM (rather than MLLM) with each question from the components combined with the description of the image and ask the LLM to score each image in the range (0, 5). We also compute the alignment of the image with the reason part of the message using CITE. Finally, we compute the PA score as the average of all the previous scores divided by 5. Fig. 3 shows an overview of our proposed method to evaluate the persuasiveness of advertisement images. (The prompts for each question are in Supp.)
3.5 Proposed Text-to-Image Approach
@@
Our evaluation results show that in contrast to current advancements in T2I generation models, these models lack creativity and persuasiveness and fail to generate images aligned with text when the input is not an explicit description. To address these issues in current existing models, we propose to first generate a description for an advertisement image $(D_{LLM})$ utilizing an LLM. We prompt the LLM with “Describe an advertisement image that conveys the following messages in detail: [$AR_{m}$]. Only return one paragraph of description without further explanation.” Next, we use the description as the prompt for the T2I method. Thus instead of prompting the T2I model with an implicit message, we use the creativity of LLMs to generate a detailed, visually explicit description, and use it as the prompt for T2I.
@@ 4 Experimental Setup
Dataset. We utilize the PittAd [18] dataset. Each image in the dataset has three to five interpretations ($AR_{m}$) that include an action component ($A_{m}$) and a reason component ($R_{m}$). Beyond the ($AR_{m}$) annotations, each image in this dataset is assigned to one or more topics. Using these topic annotations, we categorize advertisement messages into two groups: 1. Commercial Advertisements: Covering topics like Cars, Fashion, etc. 2. PSA Advertisements: Covering topics like Human Rights, Animal Rights, etc. Since annotators may select different topics for the same image, we only consider topics chosen by at least two annotators, ignoring the rest. To ensure consistency in experiments and accuracy in comparisons between commercial and PSA advertisements, we sampled 250 PSA advertisements (the total number of images that matched our criteria for PSA advertisements in the our test size of 4000) and 300 commercial advertisements from the test-set to ensure balance between the PSA and Commercial advertisements.
@@
Baseline Metrics. To the best of our knowledge, this is the first work proposing computational metrics for evaluating creativity and persuasiveness. Given the superior understanding of ads by LLMs compared to MLLMs [52, 3, 27], we use direct queries to an LLM ($C_{LLM}$ and $P_{LLM}$) as our baselines. The LLM (LLAMA3-instruct [9]) is given a description of the image from an MLLM (InternVL-V2-26B [7]). For alignment evaluation, we compare our proposed metric, CITE, with two state-of-the-art metrics from different alignment evaluation categories: VQA [25] and Image Reward [50]. Additionally, we reference a commonly used metrics in T2I generation—CLIP-score [14]. In VQA evaluation, the process begins by extracting image features using an image tokenizer. Given these features, an LLM is fine-tuned to answer the question, “Does this figure show [T2I prompt]? Answer with yes or no.” The model then outputs the probability of “yes” as the alignment score for the image. Following the paper’s recommendations, we used the CLIP-FlanT5 version of VQA. Image Reward uses human feedback on images to fine-tune its model. This evaluation approach uses the BLIP backbone to extract image features, followed by a Multi-layer Perceptron (MLP) to predict the alignment score. We further compare our proposed metric with 0-shot LLAMA3-instruct to show the importance of training the LLM.
@@
T2I Models. Since prompts generated by LLMs ($D_{LLM}$) are often longer and more detailed, the effectiveness of the text tokenizer in the T2I model influences performance when using $D_{LLM}$. After evaluating various T2I models, we selected AuraFlow-2 [11] for both $I_{LLM}$ and $I_{AR}$ due to its ability to handle long prompts effectively, and SDXL [33] was additionally chosen for $I_{AR}$ generation (but performed poorly with $I_{LLM}$ due to its text tokenizer).
@@
Human Annotation. To assess the accuracy of our proposed metrics, we selected 36 pairs of generated images (72 total) and six pairs of real images. Each pair of generated images includes one $I_{AR}$ and one $I_{LLM}$ image, in random order. Each pair was evaluated by 2 annotators, who were asked to rank the images based on the following criteria: (1) Alignment with action ($A_{m}$) and reason ($R_{m}$), (2) Persuasiveness, (3) Creativity, (4) Targeting of the correct audience, (5) Targeting correct appeal category, (6) Effective conversion of features into customer benefits, (7) Originality, (7) Imagination, (8) Elaboration, and (9) Synthesis. For 12 image pairs, we utilized Amazon Mechanical Turk, with two annotators for each pair. For the remaining 24 pairs of generated images and 6 pairs of real images, we invited 10 participants at our institution to complete Google forms, each evaluating six different pairs of images.
We report Krippendorff’s Alpha [21] to measure agreement between annotators and different metrics. To evaluate alignment agreement with the CITE score, we identify the image most frequently selected by annotators on alignment questions (alignment with $A_{m}$ and $R_{m}$) as each annotator’s chosen image. For creativity, we report the agreement between the ranking of images based on $C_{obj}$ and the annotators’ creativity rankings. For persuasiveness, we first assess the agreement between the ranking of the images based on the PA score and the annotators’ most frequently chosen image for persuasiveness across the multiple $AR_{m}$ statements used as prompts. We also analyze agreement across each question category used in calculating the PA score.
@@
Implementation Details. For LLAMA3-Instruct [9] and InternVL-2-26B [7], we used the default temperature settings with 8-bit quantization. For SDXL [33], we applied the default guidance scale, no negative prompt, and 28 inference steps. For AuraFlow-2 [11], we set a guidance scale of 5 and used 28 inference steps. All experiments were conducted on A100 GPUs. We fine-tuned LLAMA3-Instruct using the CPO trainer [49], with a batch size of 4, learning rate of 5e-5, and 3,000 training steps.
@@ 5 Results
This section evaluates the advertisement images based on their alignment with abstract messages, and their creativity and persuasiveness. We showcase the superiority of our proposed metrics—CITE, $C_{obj}$, PA—by comparing them to baseline metrics using human agreement, as well as of our strategy for text-to-image generation ($I_{LLM}$). Table 1 summarizes our results, showing our three proposed metrics (as the columns) and how images produced by different T2I methods perform under each.
@@ 5.1 Alignment
Image Commercial Ads PSA Ads
$I_{input-text}$ T2I CITE $C_{obj}$ PA CITE $C_{obj}$ PA
$I_{AR}$ SDXL 0.50 2.03 0.62 0.32 1.33 0.48
AuraFlow 0.50 2.12 0.64 0.31 1.36 0.42
$I_{LLM}$ AuraFlow 0.53 2.25 0.70 0.43 1.87 0.60
$I_{real}$ 0.55 2.28 0.98 0.49 2.04 0.60
Table 1: Comparison of text-image alignment (CITE), $C_{obj}$, and persuasiveness (PA), for real and differently generated images. The best result among generated images is bolded.
Image Commercial Ads PSA Ads
$I_{input-text}$ T2I CITE IR VQA CITE IR VQA
$I_{AR}$ SDXL 0.50 0.44 0.72 0.32 -0.05 0.75
AuraFlow 0.50 0.00 0.75 0.31 -0.21 0.72
$I_{LLM}$ AuraFlow 0.53 0.29 0.76 0.43 -0.13 0.69
$I_{real}$ 0.55 -0.69 0.75 0.49 -1.03 0.72
Table 2: Comparison of alignment of the images with text using our proposed CITE and baseline metrics. Best result per column bolded (including real images). IR = Image Reward.
Annotators Commercial PSA All
@@
H, VQA 0.38 0.24 0.31
H, CLIP-score 0.04 0.34 0.17
H, CITE (0-shot) 0.17 0.26 0.18
H, CITE (ours) 0.63 0.82 0.71
H1, H2 0.89 0.87 0.88
@@
Table 3: Agreement among human annotators (H) and alignment scores (AI metrics). H1 and H2 are the two different annotators. Highest H-AI scores bolded.
Baseline alignment metrics struggle with implicit input text. We first analyze the alignment between $I_{AR}$ and $I_{LLM}$ images generated by AuraFlow-2 [11] and SDXL [33], as well as real images, $I_{real}$. Table 2 presents baseline and proposed alignment scores for both generated and real advertisement images. Since $I_{real}$ are the ground-truth images in the dataset and $AR_{m}$ are the interpretation of real images, we expect $I_{real}$ to have higher alignment score than $I_{gen}$ (i.e. $I_{AR}$, $I_{LLM}$). Both baseline metrics (VQA and Image Reward) score the $I_{real}$ lower than the generated images, which contradicts our expectation. This highlights the limitations of these metrics when evaluating the alignment of images and implicit prompts.
@@
Table 3 shows that the agreement between human annotators in evaluation of alignment is 0.88. The agreement between human annotators and the CITE metric, Image Reward, and VQA is 0.71 (substantial agreement [28]), 0.11 (none to slight agreement [28]), and 0.31 (fair agreement [28]), resp., across all images. The gap in agreement scores between our proposed metric (CITE) and baseline metrics is even more significant for PSA advertisements. Unlike commercial ads, which often contain recognizable objects (e.g., the Gatorade in Fig. 1), PSAs typically rely on implicit messaging, as they do not promote a specific product. This larger gap in PSA evaluation demonstrates the limitations of baseline metrics in effectively assessing images generated from implicit text prompts. Fig. 4 shows examples of the alignment by baseline and proposed metrics.
@@
Figure 4: Example of images chosen by each annotator between $I_{LLM}$ and $I_{AR}$. For each pair of images, annotators select the image that better aligns with each $AR_{m}$. In each row, the value under each image indicates the score generated by the metric listed for that row. A $\checkmark$ represents the chosen image, while a $\times$ indicates the rejected image. The green circle highlights agreement with human annotations in choosing the better-aligned image, and the red circle indicates disagreement.
T2I models struggle with implicit text inputs in image generation. In Table 1 and 2 the CITE scores for generated images with $AR_{m}$ input text (i.e. $I_{AR}$) are lower than those for $I_{real}$ by 0.05 and 0.17 for commercial and PSA advertisements, respectively, indicating a performance gap in T2I models when handling implicit text inputs. Comparing $I_{AR}$ and $I_{LLM}$ in Table 2 shows that using a simple approach of first generating descriptions with an LLM for what an ad with this message should contain, and then using the description as T2I input, improves the alignment of the generated images with $AR_{m}$ by 0.03 (6%), and 0.11 (34%) on Commercial and PSA advertisements respectively.

### v2

@@ 3.4 Persuasiveness
Evaluating whether an image is persuasive involves scoring the persuasiveness techniques the image uses, as well as their effect on the intended audience. To capture the complexity of the task, we propose a persuasiveness score $P_{comp}$ which relies on multiple components. These components are well grounded in prior non-computational literature on persuasiveness [6, 37, 42, 45]. Some evaluate successful use of persuasiveness techniques, while others examine effect on the audience. We also incorporate our AIM score, resulting in $P_{comp+AIM}$. The components are:
@@
  • Audience (AU): Each advertisement is supposed to target specific groups of audiences. How well does the generated image target the audience the message targets?
@@
  • Benefit (B): One way to make the ads more persuasive is to depict the benefits of the product features for the customer. How well does the image convert the features to benefits for customers?
@@
  • Appeal Category (AP): Each advertisement is intended to target one of three appeal categories: Ethos, Pathos, or Logos. Ethos appeals to ethics and credibility; for example, when a respected celebrity endorses a product. Logos appeals to logic, using reasoning or factual information to persuade. Pathos aims to evoke an emotional response in the audience. How well does the image target the appeal category of the advertisement message?
  • Elaboration (E): How visually detailed is the image?
  • Originality (O): How out of the ordinary and unique is it?
@@
  • Imagination (I): How well does the image help the audience imagine something they never experienced?
@@
  • Synthesis (S): How well does the image connect elements that are usually unrelated?
The evaluation of the components is shown in Fig. 3. We first generate a description of the image using an MLLM. Because of the superiority of LLMs in reasoning [55, 5, 30], we prompt an LLM (rather than MLLM) with each component question. We also include as input to this LLM: (1) the description of the image, and (2) if applicable, the targeted audience and the appeal category for the ad message (detected by LLM). Then, for each “How/How well” question (Fig. 3), we ask the LLM to score each image in the range (0, 5). We compute $P_{comp}$ as the average of all the previous scores divided by 5. For $P_{comp+AIM}$, we also compute the alignment of the image with the intended reason $R_{m}$. In the supplementary file, we show the prompts for each question, and demonstrate through examples the benefit of combining the components to achieve better agreement with human scores. Combining the components acts like an ensemble model where slight inaccuracies in terms of individual components counter-balance each other. It can also be seen as accumulating evidence for persuasiveness from diverse sources for a more accurate result.
3.5 Text-to-Image Approach
@@
Our results show T2I models lack creativity and persuasiveness and fail to generate images aligned with text when the input is not an explicit description. To address these issues in existing models, we propose to first generate an expanded message for an advertisement image $(\mathcal{E}_{LLM})$ utilizing an LLM. We prompt the LLM with “Describe an advertisement image that conveys the following messages in detail: $AR_{m}$]. Only return one paragraph of description without further explanation.” Next, we use the resulting expansion as the prompt for the T2I method. Thus instead of prompting the T2I model with an implicit message, we use the creativity of LLMs to generate a detailed, visually explicit description, and use it as the prompt for T2I. We refer to the resulting images as $I_{LLM}$. In contrast, the images directly generated with the action-reason prompt are denoted $I_{AR}$.
@@ 4 Experimental Setup
Dataset. We utilize the PittAd [21] dataset. Each image has three to five interpretations ($AR_{m}$) that include an action component ($A_{m}$) and a reason component ($R_{m}$). Each image in this dataset has one or more topic annotations. We use these to categorize advertisement messages into two groups: 1. Commercial Advertisements: Covering topics like Cars, Fashion, etc. 2. PSA Advertisements: Covering topics like Human Rights, Animal Rights, etc. We only consider topics labeled in an image by at least two annotators. To ensure balance, consistency in experiments and accuracy in comparisons between commercial and PSA advertisements, we sampled 250 PSA advertisements (the total number of images that matched our criteria for PSA advertisements in the test size of 4000 from [30]) and 300 commercial advertisements from the test set.
@@
Baseline Metrics. To the best of our knowledge, this is the first work proposing computational metrics for evaluating creativity and persuasiveness. Given the superior understanding of ads by LLMs compared to MLLMs [55, 5, 30], we use direct queries to an LLM ($C_{LLM}$ and $P_{LLM}$) as our baselines. The LLM (e.g. LLAMA3-Instruct-8B [11]) is given a description of the image from an MLLM (e.g. InternVL-V2-26B [9]). For alignment evaluation, we compare our proposed metric, AIM, with three state-of-the-art text-to-image evaluation metrics: VQAScore [28], ImageReward [53], and CLIPScore [17]. In VQAScore, an LLM is fine-tuned to answer the question, “Does this figure show [T2I prompt]?” The model outputs the probability of “yes” as the alignment score for the image. Following the paper’s recommendations, we used the CLIP-FlanT5 version of VQAScore. ImageReward uses (expensive) human feedback on images to fine-tune its model. We also compare our proposed AIM metric with a zero-shot version of AIM (non-fine-tuned LLM) to show the importance of fine-tuning.
LLM Models. We test different LLMs for (1) evaluation (computing AIM), and (2) producing the expansion $(\mathcal{E}_{LLM})$. We use LLAMA3-Instruct-8B (LLAMA3) [11] and Qwen2.5-7B-Instruct (QwenLM) [2].
MLLM Models. We experiment with InternVL2-26B (InternVL) [9] and Qwen2-VL-7B-Instruct (QwenVL) [3] to produce the image description.
@@
T2I Models. Since prompts generated by LLMs ($\mathcal{E}_{LLM}$) are often longer and more detailed, the effectiveness of the text tokenizer in the T2I model influences performance. After evaluating various T2I models, we selected AuraFlow-v0.2 (AuraFlow) [13] and Flux-v1-dev (Flux) [15] for both $I_{LLM}$ (our proposed expansion method) and $I_{AR}$ (directly inputting action-reason to T2I) due to its ability to handle long prompts effectively. sdxl-flash (SDXL) [35] was additionally chosen for $I_{AR}$ generation (but performed poorly with $I_{LLM}$ due to its text tokenizer).
@@
Human Annotation. To assess the accuracy of our proposed metrics, we selected 60 pairs of generated images (120 total) and six pairs of real images. Each pair of generated images includes one $I_{AR}$ and one $I_{LLM}$ image (created with LLM expansion of the AR message), in random order. Each pair was evaluated by 2 annotators (56 unique annotators overall) asked to rank the images based on these criteria: (1) Alignment with action ($A_{m}$) and reason ($R_{m}$), (2) Persuasiveness, (3) Creativity, (4) Targeting of the correct audience, (5) Targeting correct appeal category, (6) Effective conversion of features into customer benefits, (7) Originality, (7) Imagination, (8) Elaboration, and (9) Synthesis. The annotation interface is included in the supplementary file. We report Krippendorff’s Alpha [24] to measure agreement between annotators and different metrics. To evaluate agreement with AIM, we identify the image most frequently selected by annotators, across questions on this pair about alignment with the multiple $A_{m}$ and $R_{m}$. For creativity, we report the agreement between the ranking of images based on $C_{obj}$ and the annotators’ creativity rankings. For persuasiveness, we assess the agreement between the ranking based on $P_{comp+AIM}$ and the annotators’ most frequently chosen image. We also analyze agreement across each component used to calculate persuasiveness.
@@
Implementation. We fine-tuned LLAMA3 and QwenLM using CPO [52], with a batch size of 4, learning rate of 5e-5, and 3,000 training steps. All experiments were conducted on A100 and L40s GPUs. Other details in supp.
@@ 5 Results
This section evaluates the advertisement images based on their alignment with abstract messages, and their creativity and persuasiveness. We showcase the superiority of our proposed metrics—AIM, $C_{obj}$, and $P_{comp+AIM}$—by comparing them to baseline metrics using human agreement. We also show the benefits of our strategy for text-to-image generation: expanding the message with an LLM ($I_{LLM}$).
@@ 5.1 Alignment
Annotators COM PSA All
@@
H, VQAScore 0.38 0.24 0.31
H, CLIPScore 0.04 0.34 0.17
H, AIM (InternVL, LLAMA3) (0-shot) 0.17 0.26 0.18
H, AIM (InternVL, LLAMA3) 0.60 0.82 0.68
H, AIM (InternVL, QwenLM) 0.62 0.56 0.60
H, AIM (QwenVL, LLAMA3) 0.62 0.56 0.60
H, AIM (QwenVL, QwenLM) 0.72 0.56 0.65
H1, H2 0.86 0.85 0.86
@@
Table 1: Agreement among human annotators (H) and alignment AI metrics. We show the impact of different MLLMs and LLMs for AIM (Fig. 2) by ([MLLM], [LLM]). H1, H2 are the annotators. Highest H-AI scores bolded.
Evaluation Image COM Ads PSA Ads
MLLM LLM T2I $I_{input-text}$ AIM $C_{obj}$ $P_{c+A}$ AIM $C_{obj}$ $P_{c+A}$
InternVL LLAMA3 SDXL $I_{AR}$ 0.50 2.03 0.62 0.32 1.33 0.48
AuraFlow $I_{AR}$ 0.50 2.12 0.64 0.31 1.36 0.42
$I_{LLAMA3}$ 0.53 2.25 0.70 0.43 1.87 0.60
$I_{QwenLM}$ 0.55 2.34 0.59 0.48 2.05 0.54
FLUX $I_{AR}$ 0.51 2.06 0.48 0.43 1.83 0.44
$I_{LLAMA3}$ 0.54 2.20 0.52 0.47 1.93 0.53
$I_{QwenLM}$ 0.53 2.35 0.55 0.47 2.06 0.46
QwenLM SDXL $I_{AR}$ 0.49 1.90 0.50 0.44 1.32 0.35
AuraFlow $I_{AR}$ 0.47 1.91 0.50 0.30 1.28 0.41
$I_{LLAMA3}$ 0.50 2.50 0.50 0.49 1.83 0.44
$I_{QwenLM}$ 0.51 2.18 0.50 0.48 2.04 0.45
FLUX $I_{AR}$ 0.50 1.94 0.49 0.47 1.86 0.45
$I_{LLAMA3}$ 0.51 2.04 0.49 0.48 1.93 0.45
$I_{QwenLM}$ 0.51 2.18 0.56 0.48 2.03 0.46
QwenVL LLAMA3 SDXL $I_{AR}$ 0.52 2.06 0.52 0.45 1.88 0.38
AuraFlow $I_{AR}$ 0.51 2.08 0.44 0.45 1.92 0.35
$I_{LLAMA3}$ 0.53 2.19 0.54 0.47 2.03 0.46
$I_{QwenLM}$ 0.54 2.32 0.54 0.48 2.08 0.47
FLUX $I_{AR}$ 0.51 2.02 0.47 0.46 1.89 0.42
$I_{LLAMA3}$ 0.53 2.17 0.49 0.47 1.97 0.45
$I_{QwenLM}$ 0.53 2.30 0.47 0.48 2.09 0.46
QwenLM SDXL $I_{AR}$ 0.49 1.93 0.43 0.44 1.84 0.37
AuraFlow $I_{AR}$ 0.48 1.95 0.43 0.44 1.87 0.36
$I_{LLAMA3}$ 0.50 2.18 0.46 0.46 1.97 0.43
$I_{QwenLM}$ 0.52 2.20 0.47 0.48 2.05 0.44
FLUX $I_{AR}$ 0.48 1.94 0.44 0.46 1.92 0.40
$I_{LLAMA3}$ 0.49 2.03 0.47 0.47 1.97 0.43
$I_{QwenLM}$ 0.52 2.20 0.47 0.48 2.06 0.44
Table 2: Comparison of text-image alignment (AIM), $C_{obj}$, and persuasiveness ($P_{c+A}$, short for $P_{comp+AIM}$). Best result per group is bolded. MLLM is the image descriptor. LLM is used in computing AIM and $P_{c+A}$ (and separately for input text).
@@
Baseline alignment metrics struggle with implicit input text. Table 1 - column All, which is the agreement over all the images in our human evaluation, shows that the agreement between human annotators in evaluation of alignment is 0.86. We show variants of the AIM metric using different MLLMs used for description (Fig. 2) and LLMs. The best agreement between human annotators and the AIM metric is 0.68 (substantial agreement [31]). In contrast, agreement with ImageReward is 0.11 (none to slight agreement [31]), with VQAScore is 0.31 (fair agreement [31]), and with CLIPScore is 0.17 (none to slight agreement), across all images. The gap in agreement scores between our proposed metric (AIM) and baseline metrics is even more significant for PSA advertisements (e.g. compared to VQAScore). Unlike commercial ads, which often contain recognizable objects (e.g. the Gatorade in Fig. 1), PSAs typically do not promote a specific product hence the exact content to be depicted is even more implicit. This larger gap in PSA evaluation demonstrates the limitations of baseline metrics in effectively assessing images generated from implicit text prompts. This table also demonstrates the impact of training the LLM (compared to the zero-shot version). Fig. 4 shows examples of the alignment by baseline and proposed metrics. In the supplementary, we also include results from scoring ads with baseline metrics (VQAScore, CLIPScore, ImageReward) on both real and generated images. We observe that some baseline metrics score real ad images worse than generated ones, which contradicts expectations and calls these metrics’ reliability into question.
@@
Figure 4: Example of images chosen by each annotator between $I_{AR}$ (left) and $I_{LLAMA3}$ (right). For each pair of images, annotators select the image that better aligns with each $AR_{m}$. In each row, the value under each image indicates the score generated by the metric listed. A $\checkmark$ represents the chosen (ranked better) image, while a $\times$ indicates the rejected image. The green circle highlights agreement with human annotations in choosing the better-aligned image, and the red circle indicates disagreement.
T2I models struggle with implicit text inputs in image generation. In Table 2, we compare $I_{AR}$ and $I_{LLM}$ using different choices of: (1) MLLM for image description, (2) LLMs for computing AIM, (3) text-to-image models, and (4) LLMs for message expansion (resulting in different $I_{LLM}$). The table shows that using a simple approach of first generating descriptions with an LLM for what an ad with this message should contain, and then using the description as T2I input, improves the alignment of the generated images with $AR_{m}$. Within each group of results ($I_{AR}$, $I_{LLAMA3}$ and $I_{QwenLM}$), the best AIM result is always from one of the LLMs. In most cases, $I_{QwenLM}$ outperforms $I_{LLAMA3}$. Results are not significantly different using the different MLLMs (InternVL and QwenVL) and for different evaluation LLMs (LLAMA3 and QwenLM). T2I methods AuraFlow and FLUX also have similar performance, with FLUX having an edge in some cases (e.g. PSAs using $I_{AR}$ with InternVL for MLLM).